04. 08. 2003

Language: ngerman

today  normal date format 4. August 2003
short original  4. August ’03
isodate  2003-08-04
numdate  04. 08. 2003
shortdate  04. 08. 03
TeXdate  2003/08/04
back to normal  4. August 2003
another day normal date format 30. Januar 1972
short original  30. Januar ’72
isodate  1972-01-30
numdate  30. 01. 1972
shortdate  30. 01. 72
TeXdate  1972/01/30
back to normal  30. Januar 1972
range 1  normal date format 30. bis 31. Januar 1972
short original  30. bis 31. Januar ’72
isodate  1972-01-30 bis 31
numdate  30. bis 31. 01. 1972
shortdate  30. bis 31. 01. 72
TeXdate  1972/01/30 bis 31
back to normal  30. bis 31. Januar 1972
range 2  normal date format 30. Januar bis 28. Februar 1972
short original  30. Januar bis 28. Februar ’72
isodate  1972-01-30 bis 02-28
numdate  30. 01. bis 28. 02. 1972
shortdate  30. 01. bis 28. 02. 72
TeXdate  1972/01/30 bis 02/28
back to normal  30. Januar bis 28. Februar 1972
range 3  normal date format 30. Januar 1972 bis 28. Januar 2001
short original  30. Januar ’72 bis 28. Januar ’01
isodate  1972-01-30 bis 2001-01-28
numdate  30. 01. 1972 bis 28. 01. 2001
shortdate  30. 01. 72 bis 28. 01. 01
TeXdate  1972/01/30 bis 2001/01/28
back to normal  30. Januar 1972 bis 28. Januar 2001

Language: naustrian

today  normal date format 4. August 2003
short original  4. August ’03
isodate  2003-08-04
numdate  04. 08. 2003
shortdate  04. 08. 03
TeXdate  2003/08/04
back to normal  4. August 2003
another day normal date format 30. Jänner 1972
short original  30. Jänner ’72
isodate  1972-01-30
numdate  30. 01. 1972
shortdate  30. 01. 72
TeXdate  1972/01/30
back to normal  30. Jänner 1972
range 1  normal date format 30. bis 31. Jänner 1972
short original  30. bis 31. Jänner ’72
isodate  1972-01-30 bis 31
numdate  30. bis 31. 01. 1972
shortdate  30. bis 31. 01. 72
TeXdate  1972/01/30 bis 31
back to normal  30. bis 31. Jänner 1972
range 2  normal date format 30. Jänner bis 28. Februar 1972
short original  30. Jänner bis 28. Februar ’72
isodate  1972-01-30 bis 02-28
numdate  30. 01. bis 28. 02. 1972
shortdate  30. 01. bis 28. 02. 72
TeXdate  1972/01/30 bis 02/28
back to normal  30. Jänner bis 28. Februar 1972
range 3  normal date format 30. Jänner 1972 bis 28. Jänner 2001
short original  30. Jänner ’72 bis 28. Jänner ’01
isodate  1972-01-30 bis 2001-01-28
numdate  30. 01. 1972 bis 28. 01. 2001
shortdate  30. 01. 72 bis 28. 01. 01
TeXdate  1972/01/30 bis 2001/01/28
back to normal  30. Jänner 1972 bis 28. Jänner 2001

Language: english

today  normal date format 4th August 2003
short original  4th August ’03
isodate  2003-08-04
numdate  04/08/2003
shortdate  04/08/03
TeXdate  2003/08/04
back to normal  4th August 2003
another day normal date format 30th January 1972
short original  30th January ’72
isodate  1972-01-30
numdate  30/01/1972
shortdate  30/01/72
TeXdate  1972/01/30
back to normal  30th January 1972
range 1  normal date format 30th to 31st January 1972
short original  30 to 31st January ’72
isodate  1972-01-30 to 31
numdate  30 to 31/01/1972
shortdate  30 to 31/01/72
TeXdate  1972/01/30 to 31
back to normal  30th to 31st January 1972
range 2  normal date format 30th January to 28th February 1972
short original  30th January to 28th February ’72
isodate  1972-01-30 to 02-28
numdate  30/01/ to 28/02/1972
shortdate  30/01/ to 28/02/72
TeXdate  1972/01/30 to 02/28
back to normal  30th January to 28th February 1972
range 3  normal date format 30th January 1972 to 28th January 2001
short original  30th January ’72 to 28th January ’01
isodate  1972-01-30 to 2001-01-28
numdate  30/01/1972 to 28/01/2001
shortdate  30/01/72 to 28/01/01
TeXdate  1972/01/30 to 2001/01/28
back to normal  30th January 1972 to 28th January 2001

Language: american

today  normal date format August 4, 2003
short original  August 4, ’03
isodate  2003-08-04
numdate  08/04/2003
shortdate  08/04/03
TeXdate  2003/08/04
back to normal  August 4, 2003
another day normal date format January 30, 1972
short original  January 30, ’72
isodate  1972-01-30
numdate  01/30/1972
shortdate  01/30/72
TeXdate  1972/01/30
back to normal  January 30, 1972
range 1  normal date format January 30 to 31, 1972
short original  January 30 to January 31, ’72
isodate  1972-01-30 to 31
numdate  01/30/ to 01/31/1972
shortdate  01/30/ to 01/31/72
TeXdate  1972/01/30 to 31
back to normal  January 30 to 31, 1972
range 2  normal date format January 30 to February 28, 1972
short original  January 30 to February 28, ’72
isodate  1972-01-30 to 02-28
numdate  01/30/ to 02/28/1972
shortdate  01/30/ to 02/28/72
TeXdate  1972/01/30 to 02/28
back to normal  January 30 to February 28, 1972
range 3  normal date format January 30, 1972 to January 28, 2001
short original  January 30, ’72 to January 28, ’01
isodate  1972-01-30 to 2001-01-28
numdate  01/30/1972 to 01/28/2001
shortdate  01/30/72 to 01/28/01
TeXdate  1972/01/30 to 2001/01/28
back to normal  January 30, 1972 to January 28, 2001

Language: french

today  normal date format 4 août 2003
short original  4 août 2003
isodate  4 août 2003
numdate  4 août 2003
shortdate  4 août 2003
TeXdate  4 août 2003
back to normal  4 août 2003
another day normal date format 30janvier 1972
short original  30janvier 1972
isodate  30janvier 1972
numdate  30janvier 1972
shortdate  30janvier 1972
TeXdate  30janvier 1972
back to normal  30janvier 1972
range 1  normal date format 30 au 31 janvier 1972
short original  30 au 31 janvier ’72
isodate  1972-01-30 au 31
numdate  30 au 31/01/1972
shortdate  30 au 31/01/72
TeXdate  1972/01/30 au 31
back to normal  30 au 31 janvier 1972
range 2  normal date format 30 janvier au 28 février 1972
short original  30 janvier au 28 février ’72
isodate  1972-01-30 au -02-28
numdate  30/01/ au 28/02/1972
shortdate  30/01/ au 28/02/72
TeXdate  1972/01/30 au /02/28
back to normal  30 janvier au 28 février 1972
range 3  normal date format 30 janvier 1972 au 28 janvier 2001
short original  30 janvier ’72 au 28 janvier ’01
isodate  1972-01-30 au 2001-01-28
numdate  30/01/1972 au 28/01/2001
shortdate  30/01/72 au 28/01/01
TeXdate  1972/01/30 au 2001/01/28
back to normal  30 janvier 1972 au 28 janvier 2001

Language: danish

today  normal date format 4. august 2003
short original  4. august ’03
isodate  2003-08-04
numdate  04/08/2003
shortdate  04/08/03
TeXdate  2003/08/04
back to normal  4. august 2003
another day normal date format 30. januar 1972
short original  30. januar ’72
isodate  1972-01-30
numdate  30/01/1972
shortdate  30/01/72
TeXdate  1972/01/30
back to normal  30. januar 1972
range 1  normal date format 30. til 31. januar 1972
short original  30 til 31. januar ’72
isodate  1972-01-30 til 31
numdate  30 til 31/01/1972
shortdate  30 til 31/01/72
TeXdate  1972/01/30 til 31
back to normal  30. til 31. januar 1972
range 2  normal date format 30. januar til 28. februar 1972
short original  30. januar til 28. februar ’72
isodate  1972-01-30 til -02-28
numdate  30/01/ til 28/02/1972
shortdate  30/01/ til 28/02/72
TeXdate  1972/01/30 til /02/28
back to normal  30. januar til 28. februar 1972
range 3  normal date format 30. januar 1972 til 28. januar 2001
short original  30. januar ’72 til 28. januar ’01
isodate  1972-01-30 til 2001-01-28
numdate  30/01/1972 til 28/01/2001
shortdate  30/01/72 til 28/01/01
TeXdate  1972/01/30 til 2001/01/28
back to normal  30. januar 1972 til 28. januar 2001

Language: norsk

today  normal date format 4. august 2003
short original  4. august ’03
isodate  2003-08-04
numdate  04/08/2003
shortdate  04/08/03
TeXdate  2003/08/04
back to normal  4. august 2003
another day normal date format 30. januar 1972
short original  30. januar ’72
isodate  1972-01-30
numdate  30/01/1972
shortdate  30/01/72
TeXdate  1972/01/30
back to normal  30. januar 1972
range 1  normal date format 30. til 31. januar 1972
short original  30 til 31. januar ’72
isodate  1972-01-30 til 31
numdate  30 til 31/01/1972
shortdate  30 til 31/01/72
TeXdate  1972/01/30 til 31
back to normal  30. til 31. januar 1972
range 2  normal date format 30. januar til 28. februar 1972
short original  30. januar til 28. februar ’72
isodate  1972-01-30 til -02-28
numdate  30/01/ til 28/02/1972
shortdate  30/01/ til 28/02/72
TeXdate  1972/01/30 til /02/28
back to normal  30. januar til 28. februar 1972
range 3  normal date format 30. januar 1972 til 28. januar 2001
short original  30. januar ’72 til 28. januar ’01
isodate  1972-01-30 til 2001-01-28
numdate  30/01/1972 til 28/01/2001
shortdate  30/01/72 til 28/01/01
TeXdate  1972/01/30 til 2001/01/28
back to normal  30. januar 1972 til 28. januar 2001

Changed ISO output

2000-12-24 2000–12–24 2000·12·24